the point of intersection of tangents of the circle x2 +y2 =10 drawn at the end points of the chord x+y=2 ?

Grade:11

1 Answers

Arun
6 years ago
x +y = 2.          .......(1)
 
Let this line meet circle ​S1 at A & B
Let the tangents to circle ​Sat A & B meet at P(h,k) , then AB will be chord of contact of the tangents to the circle ​Sfrom P, therefore equation of AB will be 
xh + yk - 10 = 0                   ....... (2) 

Equation (1) and (2) are same, therefore they are identical
 
1/h = 1/k = 2/10
h = 5, k = 5